OpenStudy (anonymous):

give ur answer to correct 4 significant figures. log 28+log(log 2.5)

OpenStudy (amilapsn):

are you allowed to use calculators?

OpenStudy (anonymous):

yeah,we can

OpenStudy (amilapsn):

what's the value you get?

OpenStudy (anonymous):

1.047

OpenStudy (anonymous):

but need the work

OpenStudy (amilapsn):

are there any preceding ecimals?

OpenStudy (amilapsn):

*decimals

OpenStudy (anonymous):

yeah

OpenStudy (amilapsn):

I think you should give ur answer like this: log(28)+log(log(2.5))=1.047....=1.047 in the blank you can write two or three decimals and you should round that value to 3rd decimal place right?

OpenStudy (amilapsn):

Am I correct @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\log(28)+\log(\log (2.5)) = \log(28(\log(2.5)=\log(\log(2.5^{28}))\]

OpenStudy (unklerhaukus):

yeah, i get the same 1.046975636.... = 1.047 (4 sig.figs.)

OpenStudy (amilapsn):

Thank you @UnkleRhaukus